OpenStudy (anonymous):

what is the equation for questions that give you a point and the slope and you hve to write in slope intercept form? for example: through: (4,4) slope=3/4

OpenStudy (anonymous):

(y-4)=3/4(x-4)

OpenStudy (anonymous):

I believe

OpenStudy (aroub):

You can either use the slope-intercept form: y=mx+b or the point-slope form: y-y1=m(x-x1) I like to use the slope-intercept form =)

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

each has their advantages

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

what does the point slope form find? Does it find the y int?

OpenStudy (aroub):

Yup!

OpenStudy (anonymous):

thank you!